OpenStudy (anonymous):

I need some assistance checking my work. My teacher messed up the key. :/ Write the equation of the circle with center and radius:

OpenStudy (anonymous):

\[c(1,-3), r=\sqrt{5}\]

OpenStudy (anonymous):

I got \[x^2-4x+y^2+8y+20=7\] Not sure if that is entirely right.

OpenStudy (amistre64):

(x-Cx)^2 + (y-Cy)^2 = r^2

OpenStudy (anonymous):

ER bah, I goofed

OpenStudy (anonymous):

\[C(2,-4), r \sqrt{7}\]

OpenStudy (amistre64):

x-1)^2 = x^2 -2x +1 y--3)^2 = y^2 +6y +9 r^2 = 5

OpenStudy (anonymous):

Sorry about that.

OpenStudy (amistre64):

same procedure, just new numbers :)

OpenStudy (amistre64):

x^2-4x+4 +y^2+8x+16 = 7 maybe?

OpenStudy (amistre64):

looks fine to me